Youngs Modulus (x1046 psi) 7.68 10.06 10.64 9.91 9.42 8.20 9.08 11.52 9.87 11.04 Group 2 3 4 5 6 7 10Problem 2. Using the method of least squares, find the straight line equation that represents the stress-strain relation as o

Young's Modulus (x1046 psi) 7.68 10.06 10.64 9.91 9.42 8.20 9.08 11.52 9.87 11.04 Group 2 3 4 5 6 7 10
Problem 2. Using the method of least squares, find the straight line equation that represents the stress-strain relation as obtained from your group experiment, with the strain being measured experimentally and stress computed from the theoretical analysis. Draw the curve and discuss the goodness of fit. Problem 3. Using the method of least squares, find the straight line equation that represents the values of elastic modulus as obtained from different group experiments, with the strain being measured experimentally and stress computed from the theoretical analysis. Draw the curve and discuss the goodness of fit
Answer #1

Sol Problem- 2 et us consider roup of the exp is tress let us Consider e 8tsat hat we heve to titis Accordina)右臭eart square p

Table 1
Group Youngs Modulus X 10^6 Y=8.7553+0.1794X
X Y XY X^2 Y Estimated
1 7.68 7.68 1 8.9347
2 10.06 20.12 4 9.1141
3 10.64 31.92 9 9.2935
4 9.91 39.64 16 9.4729
5 9.42 47.1 25 9.6523
6 8.2 49.2 36 9.8317
7 9.08 63.56 49 10.0111
8 11.52 92.16 64 10.1905
9 9.87 88.83 81 10.3699
10 11.04 110.4 100 10.5493
55 97.42 550.61 385
Table 2
Y=8.7553+0.1794X
Group Original Stress Estimated Stress
1 7.68 8.9347
2 10.06 9.1141
3 10.64 9.2935
4 9.91 9.4729
5 9.42 9.6523
6 8.2 9.8317
7 9.08 10.0111
8 11.52 10.1905
9 9.87 10.3699
10 11.04 10.5493

14 12 10 Original stress Estimated Stress 4 10 12

heu om the tuble 1 we have ing hese Puttny thase valanand ee equ ase A: 8.7553 요 B=0.1794 as o oup experimen Uncleaim From th